For a consumer, the marginal utility of good A is 25 and its price is $5. The marginal utility of good B is 60 and its price is $12. The consumer has allocated his entire budget. Is this consumer maximizing his total utility?
1
Expert's answer
2020-06-23T10:23:20-0400

Yes the consumer is maximizing his total utility.

according to the rule : MU1/P1 = MU2/P2

Good A Good B

25/5 = 60/12

5 = 5

Where the marginal utility per dollar is the same for both goods